### v1

Abstract
Robust adversarial reinforcement learning has emerged as an effective paradigm for training agents to handle uncertain disturbance in real environments, with critical applications in sequential decision-making domains such as autonomous driving and robotic control. Within this paradigm, agent training is typically formulated as a zero-sum Markov game between a protagonist and an adversary to enhance policy robustness. However, the trainable nature of the adversary inevitably induces non-stationarity in the learning dynamics, leading to exacerbated training instability and convergence difficulties, particularly in high-dimensional complex environments. In this paper, we propose a novel approach, Uncertainty-Aware Critic Ensemble for robust adversarial Reinforcement learning (UACER), which consists of two strategies: 1) Diversified critic ensemble: a diverse set of $K$ critic networks is exploited in parallel to stabilize Q-value estimation rather than conventional single-critic architectures for both variance reduction and robustness enhancement. 2) Time-varying Decay Uncertainty (TDU) mechanism: advancing beyond simple linear combinations, we develop a variance-derived Q-value aggregation strategy that explicitly incorporates epistemic uncertainty to dynamically regulate the exploration-exploitation trade-off while simultaneously stabilizing the training process. Comprehensive experiments across several MuJoCo control problems validate the superior effectiveness of UACER, outperforming state-of-the-art methods in terms of overall performance, stability, and efficiency.
keywords:
[…]
1 Introduction
Deep reinforcement learning (DRL) has established itself as a transformative paradigm for sequential decision-making, demonstrating remarkable success beyond gaming domains. Its applications now span real-world scenarios including precision robotic control (Huang et al., 2022; Cui et al., 2022; Valencia et al., 2024; Tang et al., 2025), adaptive healthcare systems (Yu et al., 2021a; Wang et al., 2022; Luo et al., 2024), and autonomous vehicle navigation (Feng et al., 2023; Sun et al., 2023; Zhou et al., 2024). However, a critical gap persists: most DRL policies are trained in idealized simulated environments that fail to capture the complexity of physical world deployment. This issue becomes particularly pronounced in certain high-dimensional complex tasks such as robotic control. Even minor discrepancies in physical parameters, such as friction, mass distribution, or object appearance, can lead to significant performance deterioration when policies are deployed to real-world systems. More specifically, these policies frequently overfit to simulation-specific dynamics, resulting in catastrophic performance declines when confronted with inevitable real-world disturbances (Gleave et al., 2020) or distribution shifts (Chen et al., 2021a; Zhang et al., 2024). Addressing this vulnerability, recent research (Pinto et al., 2017; Zhang et al., 2020a; Reddi et al., 2024) has prioritized the development of RL frameworks capable of counteracting adversarial disturbances and environmental variations.
Robust adversarial reinforcement learning (Pinto et al., 2017; Sheng et al., 2022; Reddi et al., 2024) is a prominent class of methods in this research domain, specifically designed to enhance the stability and generalization capabilities of RL agents operating in uncertain or adversarially perturbed environments. The framework reformulates conventional RL as a two-player zero-sum Markov game (Littman, 1994; Perolat et al., 2015), comprising a protagonist agent that optimizes policy parameters to maximize cumulative returns, and an adversary agent that systematically generates worst-case adversarial perturbations to minimize these returns. While this minimax formulation theoretically enables agents to develop resilience against worst-case perturbations, its practical deployment poses significant methodological challenges (Zhang et al., 2020b): (i) Non-Stationary Training Dynamics: The adversary’s evolving policy induces continuous perturbations, dynamically altering the protagonist’s observed state distribution and reward signals. This interaction creates a highly non-stationary learning environment where the protagonist’s optimization target shifts iteratively. As a result, policy updates exhibit high variance, frequently destabilizing the training process. (ii) Convergence Impediments: The cyclical adversarial training regime forces the protagonist to adapt to an ever-changing adversary policy before its own policy converges. Without a fixed optimization target, the protagonist remains trapped in a persistent reactive adaptation loop, leading to significantly prolonged convergence times due to continual policy adjustments, or even completed convergence failure in cases of extreme non-stationary conditions.
To address the aforementioned stability and convergence problems, we propose Uncertainty-Aware Critic Ensemble robust adversarial Reinforcement learning (UACER), a simple yet effective general framework that advances conventional robust adversarial RL by systematically integrating ensemble methodology with uncertainty quantification. This integration significantly improves both training stability and convergence speed in adversarial environments. The technical foundation of UACER addresses a fundamental limitation of single-critic architectures, which are particularly vulnerable to adversarial perturbations and distributional outliers, leading to unstable Q-value estimation (An et al., 2021a; Peer et al., 2021). To overcome this limitation, UACER employs a parallel ensemble of critic networks that provides three key benefits: Firstly, aggregated Q-value reduces the variance in the Q-function estimate and policy gradient updates so as to enhance training stability. Secondly, it effectively mitigates training oscillations and increases resilience against adversarial disturbances through diversified value function approximations. Thirdly, the increased stability and robustness facilitate accelerated policy convergence during iterative adversarial interactions. For the final Q-value determination session, we present a Time-varying Decay Uncertainty (TDU) Q-value aggregation function for combining outputs across multiple critic networks. The TDU mechanism is designed to: (i) promote extensive exploration during initial training phases through higher uncertainty tolerance, and (ii) gradually shift towards more conservative value estimation in later stages to ensure robust convergence. This way, we can enable effective balance between exploration and exploitation together with maintaining resilience against adversarial perturbations throughout the learning process. Our contributions can be summarized as follows:
  • 1)
    We develop a novel uncertainty-aware critic ensemble method for robust adversarial reinforcement learning (called UACER), which achieves more stable and accurate Q-value estimation through employing multiple base critic predictors.
  • 2)
    We design a TDU function, a principled decision strategy to combine Q-value predictions from multiple critics. This mechanism dynamically balances exploration and exploitation while maintaining stable policy updates throughout the training process.
  • 3)
    Quantitative experimental results on several complex MuJoCo problems demonstrate that UACER performs consistently better than state-of-the-art methods, including QARL and related baselines, in terms of overall performance, stability, and efficiency.
2 Related Work
Robust Adversarial Reinforcement Learning (RARL).
The concept of adversarial training for robust RL was firstly established by RARL (Pinto et al., 2017), which formulated the problem as a zero-sum Markov game between a protagonist agent and an adversarial opponent. This paradigm demonstrated that exposing agents to strategically generated perturbations during training could significantly enhance their robustness against real-world distributional shifts. Building upon this framework, subsequent research has developed specialized improvements targeting several key limitations. For instance, RARARL (Pan et al., 2019) incorporated explicit risk modeling to reduce catastrophic failures, while NR-MDP (Tessler et al., 2019) expanded the adversary’s action space to generate more diverse and challenging disturbances. Further optimization advances introduced subsequently by MixedNE-LD (Kamalaruban et al., 2020) and Cen et al. (2021) employed stochastic gradient Langevin dynamics and extragradient methods, respectively, to improve the robustness of adversarial training. More recently, A2P (Liu et al., 2024) proposed dynamically adjusting the intensity of action-space perturbations based on the agent’s current relative performance, enabling more adaptive adversarial training. While these approaches have progressively addressed specific robustness challenges, they remain limited in handling the fundamental stability and convergence inherent in adversarial training. Our work advances this line of research by developing a unified framework that simultaneously improves training stability and convergence properties, while preserving and further enhancing the robustness benefits of adversarial training paradigms.
Training Stability and Convergence Improvements for RARL.
While RARL has markedly advanced the robustness of RL policies, its adversarial formulation introduces fundamental challenges in training stability and convergence. These limitations are not confined to complex settings but are in inherent to the framework itself, as rigorously demonstrated by Zhang et al. (2020b), who showed their persistence even in simple linear quadratic regulation problems. In response, recent research has primarily evolved along three directions: (i) Constrained optimization approaches, where DI-CARL (Zhai et al., 2022) established stability guarantees through dissipation inequation constraint and $L_{2}$ gain performance in adversary RL; Yu et al. (2021b) proposed an accelerated convergence method for linear quadratic games by employing Bregman divergence to effectively capture and adapt to the global structure of the constraint set. (ii) Curriculum learning methods enhance stability and robustness by progressively exposing the agent to escalating adversarial difficulty, notably through CAT (Sheng et al., 2022), which employed a structured curriculum to systematically increases task complexity, thereby ensuring stable and efficient training process. (iii) hybrid frameworks such as QARL (Reddi et al., 2024) that combine entropy regularization with quantal response equilibrium and use a temperature parameter to systematically manage adversary rationality in a curriculum learning manner, thereby accelerating convergence while easing the complexity of the saddle point optimization problem. In contrast, our work adopts an ensemble-based perspective that provides inherent stability to adversarial RL architecture through diversified Q-value estimation. This approach maintains compatibility with existing optimization strategies and further enhances the stability and convergence properties.
Q-Function Ensemble in Reinforcement Learning.
The integration of multiple Q-value estimators, known as Q-function ensemble, has become a widely adopted approach to enhance the stability and robustness of value-based deep RL. Early ensemble methods in the DQN family primarily address overestimation bias through averaged or pessimistic selection strategies, such as the simple averaging in Averaged DQN (Anschel et al., 2017) and the min-Q operator used in TD3 (Fujimoto et al., 2018), SAC (Haarnoja et al., 2018), and Maxmin Q-learning (Lan et al., 2020). This conservative estimation approach was further extended by methods like MSG (Ghasemipour et al., 2022) and EDAC (An et al., 2021b), which employed lower confidence bounds to systematically reduce overestimation. Beyond bias reduction, Q-function ensemble also facilitates exploration. Bootstrapped DQN (Osband et al., 2016) promoted behavioral diversity via bootstrap sampling and multi-headed Q-networks, while UCB Exploration (Chen et al., 2017) aggregated mean and variance from $K$ critics to guide action selection with an upper confidence bound. Similarly, UA-DDPG (Kanazawa et al., 2022) supported flexible ensemble sizes and encouraged exploration by prioritizing actions with high epistemic uncertainty. To further mitigate instability and balance exploration-exploitation, SUNRISE (Lee et al., 2021) was proposed as a unified ensemble RL framework, which re-weighted target Q-values based on ensemble uncertainty selects actions via UCB. Despite empirical success in standard RL settings, these methods are primarily designed for stationary, non-adversarial environments, leaving their application in robust adversarial RL largely unexplored. Under adversarial disturbances, Q-value estimates exhibit significant fluctuation, and simplistic pessimistic value estimation can hinder policy learning. Unlike prior work, we introduce a diversified critic ensemble to stabilize Q-learning against strategic perturbations and propose a time-varying decay uncertainty mechanism for Q-value aggregation, tailored to promote rapid exploration and policy learning in non-stationary adversarial RL training.
3 Problem Formulation
We consider a robust adversarial RL framework modeled as a two-player $\gamma$ discounted zero-sum Markov game (Littman, 1994; Perolat et al., 2015): $\mathcal{M}=\langle\mathcal{S},\mathcal{A}_{p},\mathcal{A}_{a},\mathcal{P},\mathcal{R},\\ \gamma\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}_{p}$ and $\mathcal{A}_{a}$ are the action spaces of the protagonist and adversary, respectively, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}_{p}\times\mathcal{A}_{a}\times\mathcal{S}\rightarrow\mathbb{R}$ defines the transition probability density, $\mathcal{R}:\mathcal{S}\times\mathcal{A}_{p}\times\mathcal{A}_{a}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward of both players, and $\gamma\in[0,1)$ is the discount factor. Our goal is to learn the policy of the protagonist (denoted by $\pi_{p}$) such that it can obtain maximum return in adversarial environments. Supposing the policy of the adversary is denoted by $\pi_{a}$, the optimization objective of the protagonist is to seek to maximize the following expected return:
$J_{\pi_{p},\pi_{a}}(s)=\mathbb{E}_{a_{p}\sim\pi_{p}(\cdot|s),a_{a}\sim\pi_{a}(\cdot|s)}\!\left[\sum\limits_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{p}^{t},a_{a}^{t},s_{t+1})\right],$
while the adversary simultaneously attempts to minimize it through its policy $\pi_{a}$. The optimal solution corresponds to a minimax or Nash equilibrium satisfying:
$J_{\pi_{p}^{*},\pi_{a}^{*}}(s)=\min_{\pi_{a}}\max_{\pi_{p}}J_{\pi_{p},\pi_{a}}=\max_{\pi_{p}}\min_{\pi_{a}}J_{\pi_{p},\pi_{a}}.$ (2)
In the practical implementation, the adversarial training follows an alternating procedure: at each iteration, we first hold the protagonist’s policy fixed and update the adversary’s policy to minimize the expected return $J$, then conversely freeze the adversary’s policy and learn the protagonist’s policy to maximize the return. This procedure introduces training instability and convergence difficulties due to concurrent policy updates and adversarial perturbations. Our work specifically addresses these challenges through ensemble-based value estimation and adaptive uncertainty-aware policy optimization.
4 Methodology
In this section, we give a detailed description of UACER whose overview is illustrated in Figure 1. We first introduce the diversified critic ensemble framework for robust adversarial RL and then develop the variance-based TDU Q-value aggregation function, followed by theoretical guarantees for TDU’s convergence properties and aggregation effectiveness. In principle, our method can be used in conjunction with most modern robust adversarial RL algorithms, including recent advances like QARL (Reddi et al., 2024) as well as classical approaches such as RARL (Pinto et al., 2017). For the sake of concreteness, we here provide the QARL-based implementation as our primary instantiation.
Figure 1: Schematic of the proposed UACER method.
4.1 Diversified Critic Ensemble
The Unified Framework.
The UACER framework implements both protagonist and adversary as actor-critic agents based on the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018). Each agent consists of a single policy network, denoted as $\pi_{\phi}$, and an ensemble of $K$ soft Q-function networks $\{Q_{\theta_{k}}\}_{k=1}^{K}$. This design is motivated by the empirical insight, as shown in Figure LABEL:actor_ensemble, that the critic ensemble significantly enhances the robustness of Q-value estimation, especially under adversarial perturbations. By evaluating Q-values through multiple independent critics, the framework mitigates the influence of localized noise or perturbations on any single estimator, thereby yielding more stable and reliable value estimates. In contrast, although actor ensembles can improve sample efficiency and generalization (Lee et al., 2021; Yang et al., 2022), they often introduce additional training instability due to policy divergence among actors, which may produce conflicting behaviors that undermine the consistency and convergence of the learned policy.
During the adversarial training phase, both agents interact synchronously with the environment through their current policies, generating transition tuples $\tau_{t}=(s_{t},a_{p_{t}},a_{a_{t}},r_{t},\\ s_{t+1})$ that are stored in the experience replay buffer $\mathcal{D}$. Given the symmetric nature of the protagonist-adversary learning paradigm, their parameter update principle can be uniformly formalized as follows: At timestep $t$, the protagonist executes action $a_{t}=a_{p_{t}}$ and receives reward $R_{t}=r_{t}$, while the adversary selects action $a_{t}=a_{a_{t}}$ and obtains reward $R_{t}=-r_{t}$. To reduce the variance in the Q-function estimate, we update parameter $\theta_{k}$ for each critic $Q_{\theta_{k}}(k\in\{1,2,\dots,K\})$ with the same target to minimize the soft Bellman residual:
$\mathcal{L}_{Q}(\theta_{k})=\mathbb{E}_{\tau_{t}\sim\mathcal{D}}\Big[\big(Q_{\theta_{k}}(s_{t},a_{t})-R_{t}-\gamma V(s_{t+1})\big)^{2}\Big],$ (3)
with
$V(s_{t})=\mathbb{E}_{a_{t}\sim\pi_{\phi}}\big[Q_{\bar{E}}(s_{t},a_{t})-\alpha\log\pi_{\phi}(a_{t}|s_{t})\big],$ (4)
where $\alpha$ is the temperature parameter, and $Q_{\bar{E}}(s_{t},a_{t})$ represents the ensemble Q-value estimate aggregated from $K$ target critics through our proposed TDU function, whose formal definition will be detailed in the next section. Then, the policy $\pi_{\phi}$ can be updated by ensemble Q-value estimate and minimizing:
$\mathcal{L}_{\pi}(\phi)=\mathbb{E}_{s_{t}\sim\mathcal{D}}\Big[\mathbb{E}_{a_{t}\sim\pi_{\phi}}\big[\alpha\log\pi(a_{t}|s_{t})-Q_{E}(s_{t},a_{t})\big]\Big],$ (5)
where $Q_{E}(s_{t},a_{t})$ is the TDU output aggregated from $K$ critics.
(a) Robustness evaluation curves under different methodological configurations.
Diversity Enhancement.
To encourage sufficient diversity among critics, we employ a dual randomization strategy combining parameter initialization and architectural design. Each critic network undergoes independent Gaussian weight initialization with parameter noise injection, establishing distinct initial optimization paths (Plappert et al., 2018). Additionally, we randomly assign activation functions (ReLU, LeakyReLU, ELU) (Dubey et al., 2022) to each layer during forward propagation. This combined parametric and structural randomization systematically promotes representational diversity across critics while maintaining functional compatibility, with experimental results in Figure LABEL:diversity empirically validating its effectiveness.
4.2 Time-varying Decay Uncertainty Mechanism
Conventional ensemble methods in RL typically stabilize training by averaging critic outputs, which has been proven effective in reducing variance of target approximation errors (Anschel et al., 2017) and refining policy optimization directions (Chen et al., 2021b). Building on this foundation, we propose the TDU function, formally defined in Definition 1, as an optimistic aggregation mechanism tailored for $K$-critic ensembles in non-stationary environments. The TDU mechanism explicitly considers epistemic uncertainty among critics through variance quantification and dynamically modulating its impact on Q-value estimation via time-decaying coefficients. This dual mechanism design enables TDU to simultaneously enhance Q-value estimation stability and maintain an adaptive exploration-exploitation balance throughout the training process.
Definition 1 (TDU Function)
Consider a set of $K$ independent and identically distributed (i.i.d.) unbiased estimators $\{Q_{\theta_{k}}(s_{t},a_{t})\}_{k=1}^{K}$ for the true state-action value function $Q^{\ast}(s_{t},a_{t})$, where each estimator has finite variance. For the current adversarial training alternating iteration index $n\in\{1,2,...,N\}$ (where $N$ is the total number of iterations), the TDU function is defined as:
$Q_{E}(s_{t},a_{t})=\mu_{Q}(s_{t},a_{t})+\beta(n)\cdot\sigma_{Q}(s_{t},a_{t}),$ (6)
with
$\mu_{Q}(s_{t},a_{t})=\frac{1}{K}\sum_{k=1}^{K}Q_{\theta_{k}}(s_{t},a_{t}),$ (7)
$\sigma_{Q}^{2}(s_{t},a_{t})=\frac{1}{K-1}\sum\limits_{k=1}^{K}\big[Q_{\theta_{k}}(s_{t},a_{t})-\mu_{Q}(s_{t},a_{t})\big]^{2},$ (8)
and
$\beta(n)=\beta_{0}\cdot e^{-\lambda n/N}+\beta_{\min},\quad\text{s.t.}\quad\beta_{0}+\beta_{\min}=1$ (9)
is the coefficient to control the degree of considered uncertainty with time-varying decay, where $\lambda$ is a large positive constant, $\beta_{0}$ is the initial coefficient value, and $\beta_{\min}$ is a small positive value approaching zero.
As evident from the above definition, the efficacy of the TDU function is principally governed by its time-varying decay coefficient $\beta(n)$, which dynamically modulates how epistemic uncertainty contributes to Q-value aggregation. Figure LABEL:coefficient compares the robustness performance of our proposed natural exponential decay formulation for $\beta(n)$ (given in Eq. (9)) with three baseline uncertainty-weighting strategies on the MuJoCo Hopper-Hop task: (i) uncertainty-agnostic: $\beta(n)\equiv 0$, equivalent to standard averaging; (ii) constant optimistic weighting: $\beta(n)\equiv 1$, fixed maximum uncertainty weighting; (iii) linear decay: $\beta(n)$ decreasing linearly from $1$ to 0. The results demonstrate that time-decaying weighting strategies enable agents to learn significantly more robust policies compared with both uncertainty-agnostic and constant optimistic weighting approaches. This advantage stems from their ability to automatically transition from exploration-emphasizing behavior in early training stages to confidence-prioritizing predictions in later phases, thereby improving both estimation accuracy and learning efficiency. Furthermore, our proposed natural exponential decay strategy achieves superior robustness performance and faster convergence than linear decay through its monotonically decreasing derivative, allowing faster reduction of the agent’s uncertainty preference during early training while preserving essential exploration capacity for effective learning progression.
4.3 Theoretical Guarantees for TDU
In this section, we theoretically analyze the TDU mechanism through Theorem 1, establishing its convergence guarantees as a Q-value estimator in adversarial RL settings.
Theorem 1 (Unbiasedness and Consistency)
Suppose there exist $K$ critics $\{Q_{\theta_{k}}(s_{t},a_{t})\}_{k=1}^{K}$ designed to estimate the true state-action value function $Q^{\ast}(s_{t},a_{t})$, each satisfying:
  • unbiased estimation:
    $E\big[Q_{\theta_{k}}(s_{t},a_{t})\big]=Q^{\ast}(s_{t},a_{t}),\forall k\in\{1,\dots,K\};$ (10)
  • bounded variance:
    $\operatorname{Var}\big[Q_{\theta_{k}}(s_{t},a_{t})\big]\leq\sigma^{2},\forall k\in\{1,\dots,K\}.$ (11)
Then,
$\lim_{N\rightarrow\infty}\mathbb{E}\bigg[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\bigg]\leq\beta_{\min}\cdot\sigma.$ (12)
As $\beta_{\min}\to 0$, the ensemble estimator $Q_{E}(s_{t},a_{t})$ converges in probability to $Q^{*}(s_{t},a_{t})$:
$\hskip 45.00006ptQ_{E}(s_{t},a_{t})\overset{p}{\to}Q^{\ast}(s_{t},a_{t}).$ (13)
UACER is built upon established robust adversarial RL frameworks, including RARL (Pinto et al., 2017), MixedNE-LD (Kamalaruban et al., 2020), and QARL (Reddi et al., 2024), and thus inherits their theoretical guarantee of convergence to a Nash equilibrium under adversarial interactions. This foundation ensures stable and well-founded learning dynamics even in the presence of perturbations and strategic opponents. Moreover, Theorem 1 formally establishes the asymptotic reliability of TDU’s Q-value estimation. It shows that as the number of training iterations grows sufficiently large and $\beta_{\min}$ approaches zero, the expected bias between TDU’s Q-value estimates and the true Q-values diminishes asymptotically. This result confirms TDU’s consistency as an aggregation method and demonstrates its theoretical capability to converge to the optimal value function given sufficient training. The detailed proof of Theorem 1 is provided in Appendix A.
5 Experimental Results
To comprehensively evaluate our method, we conduct experiments centered around the following research questions:
  • [leftmargin=2em]
  • Q1
    Can UACER successfully achieve more robust (see Figures 4, 5 and Table 2), stable (see Table 1), and efficient (see Figure 4) adversarial RL compared with existing methods?
  • Q2
    How crucial is each proposed component in UACER for improving RL performance (see Figure 6)?
  • Q3
    Does the optimistic design of TDU provide measurable benefits in robustness and convergence compared to conventional pessimistic value estimation under adversarial conditions (see Figure 7)?
  • Q4
    How does critic ensemble size $K$ affect the performance (see Figure 9)?
  • Q5
    Can UACER be universally applicable to other existing robust adversarial RL algorithms (see Figure 8)?
Environments.
We evaluate our proposed method on a diverse set of MuJoCo control problems (Todorov et al., 2012) from the DeepMind Control Suite (Tunyasuvunakool et al., 2020), including Cartpole-Swingup Sparse, Walker-Run, Hopper-Hop, and Cheetah-Run. Following established adversarial training protocols (Pinto et al., 2017; Kamalaruban et al., 2020; Reddi et al., 2024), we instantiate two agents: a protagonist, which learns an optimal policy to accomplish the task, and an adversary, which exerts adversarial forces to disrupt the protagonist’s performance, thereby eliciting robust agent behavior. Both agents observe the standard environmental states as inputs. The adversary’s action space consists of bounded, multi-dimensional disturbance forces. For each environment, the adversary’s perturbation intensity is carefully tuned so as to be large enough to beget agent robustness and generalization while posing a challenge to the protagonist. Additional environment-specific parameter details are provided in Appendix B.1.
Comparison Methods.
We evaluate our approach in comparison to the classical robust adversarial RL method RARL (Pinto et al., 2017) and several recently proposed improvements addressing RARL’s saddle point optimization challenges, including MixedNE-LD (Kamalaruban et al., 2020), which utilizes Langevin dynamics to escape local optima, and Curriculum Adversarial Training (CAT) (Sheng et al., 2022) featuring dynamic adversary strength adjustment via a hand-designed curriculum. Additionally, we compare against the state-of-the-art QARL algorithm (Reddi et al., 2024), which synergistically combines entropy regularization with quantal response equilibrium while employing temperature-based curriculum learning to regulate adversary rationality. For comprehensive assessment, standard SAC (Haarnoja et al., 2018) serves as a non-adversarial baseline representing the underlying RL framework. Unless specified otherwise, UACER implementations in the following build upon the QARL framework. Complete hyperparameter configurations for all algorithms are provided in Appendix B.2.
Evaluation Protocols.
For each comparison method, we conduct training over five random seeds $\{0,1,2,3,4\}$ across all environments, with $200$ alternating iterations per seed. All reported results represent the average performance over these five runs. In the performance curves, each line depicts the average return obtained for varying properties of the environment after the corresponding training iteration. Solid lines indicate the mean performance, and shaded regions correspond to the $95\%$ confidence intervals across the five seeds. To evaluate the disturbance resistance of the trained protagonists to varying test conditions during training, we perform robustness evaluation every five alternating iterations. Specifically, for each environment, we assess the protagonist’s returns by sweeping through a predefined range of environmental parameter variations (e.g., mass and friction coefficient) in the absence of the adversary, and report the average return across all parameter configurations as the robustness metric. This procedure introduces graded disturbances through varying dynamics parameters, thereby offering a comprehensive test of the agent robustness while simultaneously assessing its generalization capability under domain shift conditions. Concurrently, we quantify training stability by computing the percentage decrease in the protagonist’s performance between consecutive (i.e., $i$ and $i+1$) robustness evaluations, then averaging these percentage decreases throughout the entire training process. In addition, following the practice in (Reddi et al., 2024), we assess the final adversarial performance of the protagonist in a minimax sense by fixing the protagonist obtained at the end of training and subsequently training an adversary against it, which is then used at test time.
Figure 4: Robustness evaluation curves of UACER and baseline methods on MuJoCo swingup and locomotion tasks.
Figure 5: Heatmap of final robustness evaluation for UACER and baseline methods on MuJoCo swingup and locomotion tasks.
5.1 Comparative Evaluation (Q1)
Robustness and Convergence.
As shown by the robustness evaluation results in Figure 4, UACER exhibits substantially accelerated convergence relative to all baseline methods across most environments, effectively addressing the persistent challenge of slow convergence in robust adversarial RL. This framework consistently achieves superior robustness scores across all environments, demonstrating both exceptional adversarial resilience and generalization capacity. Particularly in complex (Hopper-Hop) and sparse reward (Cartpole-Swingup Sparse) settings, UACER’s time-varying decay uncertainty-aware mechanism supports more efficient exploration and delivers statistically significant performance gains over competing approaches. Moreover, we visualize the robustness of the trained protagonist to different test conditions upon training completion, as illustrated in Figure 5. Each heatmap depicts the performance obtained under varying environmental parameters represented along the x- and y-xes. These results further validate the comprehensive robustness advantage of UACER under graded environmental perturbations.
Method Cartpole-Swingup Sparse Walker-Run Hopper-Hop Cheetah-Run
SAC $60.504\pm 53.811$ $353\pm 1.862 33.$$3.367\pm 8.704$ $4.520\pm 4.724$
RARL $56.423\pm 40.918$ $3.935\pm 2.105$ $30.215\pm 11.733$ $6.943\pm 2.906$
SAC MixedNE-LD $88.317\pm 4.567$ $3.151\pm 1.376$ $36.220\pm 14.331$ $6.281\pm 2.281$
CAT $56.902\pm 51.160$ $4.743\pm 2.955$ $32.276\pm 12.829$ $5.353\pm 3.882$
QARL $35.714\pm 28.414$ $4.527\pm 0.847$ $35.581\pm 6.025$ $8.692\pm 4.927$
UACER (ours) $17.352\pm 9.593$ $1.659\pm 0.604$ $27.938\pm 8.474$ $7.665\pm 3.515$
Table 1: Training stability analysis (percentage decrease, $\%$) of UACER and baseline methods on MuJoCo swingup and locomotion tasks.
Method Cartpole-Swingup Sparse Walker-Run Hopper-Hop Cheetah-Run
RARL $107.420\pm 182.710$ $257.679\pm 50.298$ $14.267\pm 15.725$ $268.101\pm 52.009$
SAC MixedNE-LD $0.000\pm 0.000$ $233.273\pm 41.541$ $15.222\pm 14.110$ $227.447\pm 54.054$
CAT $0.000\pm 0.000$ $229.088\pm 23.850$ $9.514\pm 10.039$ $192.526\pm 14.081$
QARL $111.420\pm 189.457$ $253.536\pm 46.556$ $9.386\pm 8.461$ $292.861\pm 42.095$
UACER (ours) $334.770\pm 6.886$ $341.995\pm 8.480$ $96.564\pm 17.479$ $312.594\pm 30.422$
Table 2: Final adversarial performance of UACER and baseline methods on MuJoCo swingup and locomotion tasks.
Figure 6: Robustness evaluation curves of UACER and its three variants (UACER without ensemble/TDU/diversity) on MuJoCo swingup and locomotion tasks.
Training Stability.
We evaluate the training stability of UACER from two aspects: variability across different random seeds and consistency of returns during training. Figure 4 shows UACER achieves the smallest variance across seeds (narrowest shaded regions), indicating the lowest sensitivity to initialization for training. Furthermore, training stability metrics derived from the results in Figure 4 are summarized in Table 1, where performance (percentage decrease, $\%$) is reported as mean $\pm$ standard error across five independent trials, with bold and underlined values denoting the best and second-best results, respectively. The results show that UACER achieves the best training stability in the first three environments while maintaining comparable performance to the baselines in the last environment (i.e., Cheetah-Run). These results further validate its effectiveness in mitigating adversarial-induced training fluctuations in robust adversarial RL settings.
Worst-Case Performance.
As quantified final adversarial performance in Table 2, UACER consistently maintains superior performance with a relatively small variance across all environments when confronting optimized adversaries. This validates its dual capability of preserving policy effectiveness under distributional shifts while enhancing generalizability in extreme out-of-distribution scenarios.
5.2 Ablation Study (Q2)
We systematically evaluate the individual contributions of UACER’s core components: the critic ensemble, the TDU aggregation mechanism, and the diversity enhancement strategies. Through controlled single-variable ablation experiments, we compare three algorithmic variants in adversarial RL settings: (i) w/o ensemble: Replaces the $K$-critic ensemble with SAC’s standard two-critic architecture; (ii) w/o TDU: Substitutes TDU aggregation function with minimum-value selection used in SAC; (iii) w/o diversity: Removes the diversity enhancement strategies from UACER.
The experimental results in Figure 6 demonstrate that removing any component from UACER degrades performance substantially. When eliminating the diversity enhancement strategies ("w/o diversity"), the algorithm exhibits markedly poorer training stability and robustness across most environments. More critically, the removal of either the TDU mechanism ("w/o TDU") or the critic ensemble ("w/o ensemble") leads to notable performance declines in both robustness and convergence speed, particularly in challenging environments such as Cartpole-Swingup Sparse and Hopper-Hop. In contrast, the fully integrated UACER framework consistently outperforms all ablated versions. These results indicate that under highly perturbed environments, no individual component suffices for robust performance; instead, their synergistic integration is essential.
5.3 Superiority of TDU’s Optimistic Aggregation in Adversarial RL (Q3)
[…]
To evaluate the superiority of TDU’s optimistic aggregation in adversarial RL, we compare it with three pessimistic variants derived from Eq. (6): (i) pessimism-dec: the mean minus the uncertainty term, with $\beta(n)$ decaying from 1 to 0; (ii) pessimism-inc: the mean minus the uncertainty term, with $\beta(n)$ increasing from 0 to 1; (iii) pessimism-min: the minimum among the $K$ critic outputs. As demonstrated in Figure 7, TDU consistently delivers superior performance across all tasks. This advantage arises because, in non-stationary adversarial environments, the protagonist already struggles to learn appropriate responses to perturbations, and pessimistic estimation further constraints its Q-value prediction capacity under such disturbances. By contrast, TDU’s optimistic aggregation achieves a more effective balance between bias control and robustness enhancement, thereby alleviating the tendency toward over-conservative policy learning while simultaneously strengthening training stability and final performance.
Figure 7: Robustness evaluation curves of UACER (optimistic TDU aggregation) and its three pessimistic Q-value estimation variants on MuJoCo swingup and locomotion tasks.
Figure 8: Robustness evaluation curves of UACER combined with RARL/SAC MixedNE-LD on MuJoCo swingup and locomotion tasks.
Figure 9: Robustness evaluation of UACER with varying values of critic ensemble size $K$ on the Hopper-Hop environment.
5.4 Hyperparameter and Scalability Analysis
Effects of Critic Ensemble Size (Q4)
The ensemble size $K$ (number of critics) in UACER critically influences the agent’s performance. To investigating this relationship, we conduct controlled experiments varying $K$ from $2$ to $10$ while holding other hyperparameters fixed. As illustrated in Figure 9, we observe that the final obtained return in terms of robustness evaluation improves gradually with increasing $K$ up to $K=5$, after which performance declines. This indicates that while larger ensembles enhance uncertainty estimation and ensemble diversity, excessive critics ($K>5$) may introduce detrimental redundancy, complicating policy optimization. Consequently, a higher value of $K$ is not always better. Here, for the Hopper-Hop task, $K=5$ achieves the optimal robustness gains, suggesting task-independent determination of $K$ is essential.
Scalability Analysis (Q5)
All previous experiments implemented UACER based on the QARL framework as the foundational robust adversarial RL approach. This section examines UACER’s applicability as a general framework when integrated with other existing robust adversarial RL algorithms. Specifically, we migrate UACER to the classic RARL algorithm and its improved extension SAC MixedNE-LD for practical implementation and evaluation. The learning curves in terms of robustness performance are shown in Figure 8. The experimental results confirm that both RARL and SAC MixedNE-LD achieve substantial performance gains after incorporating UACER, showing improved convergence speed, training stability, and robustness to adversarial perturbations. In essence, UACER is designed as an algorithm-agnostic framework that is compatible with any value-based robust adversarial RL methods. Regrettably, in the Cartpole-Swingup Sparse task, SAC MixedNE-LD exhibited critically poor baseline performance (near-zero returns throughout training), with no discernible improvement after UACER integration. This outcome indicates that UACER’s performance improvements remain partly dependent on the base adversarial RL algorithm’s intrinsic adaptability to the specific task.
5.5 Computational Overhead and Limitations
A potential limitation of UACER is the additional computational cost introduced by updating multiple critics. Nevertheless, this overhead can be effectively mitigated through parallel execution, which prevents wall-clock training time from rising substantially compared to single-critic baselines. By performing critic updates concurrently, the incremental per-iteration computation is effectively absorbed, maintaining overall training efficiency. In return, UACER delivers notable improvements in training stability, convergence speed, and robustness against adversarial perturbations. Given these performance gains, the elevated computational demand remains acceptable and manageable in practice, representing a favorable trade-off relative to conventional single-critic adversarial RL architectures.
6 Conclusion
We propose UACER, a novel approach that significantly advances robust adversarial RL methods through two key contributions: (i) a parallel ensemble of critic networks that simultaneously reduces Q-value estimation variance and enhances robustness against adversarial perturbations, and (ii) a Time-varying Decay Uncertainty (TDU) Q-value aggregation mechanism that dynamically balances exploration and exploitation through variance-based epistemic uncertainty quantification across critics. Theoretically grounded and extensively validated, UACER demonstrates substantial improvements in training stability, convergence speed, and adversarial robustness, addressing critical challenges for real-world deployment of deep RL systems. By integrating ensemble learning with uncertainty-aware optimization, our framework enables the development of adversarial RL algorithms with proven stability guarantees and robust performance in complex and uncertain environments. Future research directions include extending UACER to partially observable environments and investing its applications in real-world safety-critical systems such as robotic manipulation and autonomous driving scenarios.
Appendix A Proof of Theorem 1
In this section, we present the proof of Theorem 1. Using the conditions specified in Definition 1, we establish the theorem’s validity as follows.
Proof 1
Given the unbiasedness condition $E[Q_{\theta_{k}}(s_{t},a_{t})]=Q^{\ast}(s_{t},a_{t})$ for all $k\in\{1,\dots,K\}$, it follows that the ensemble mean satisfies
$\mathbb{E}\big[Q_{\theta_{k}}(s_{t},a_{t})\big]=\mu_{Q}(s_{t},a_{t})=Q^{*}(s_{t},a_{t}).$ (14)
Consider the deviation of the ensemble estimate $Q_{E}(s_{t},a_{t})$ from the true value $Q^{\ast}(s_{t},a_{t})$:
$\big|Q_{E}$ $(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|$ (15)
$=\big|\mu_{Q}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})+\beta(n)\cdot\sigma_{Q}(s_{t},a_{t})\big|.$
Then, applying the triangle inequality yields
$\big|Q_{E}$ $(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|$ (16)
$\leq\big|\mu_{Q}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|+\beta(n)\cdot\sigma_{Q}(s_{t},a_{t}).$
Taking expectations on both sides and let the number of training iterations $N\to\infty$, we obtain
[…]
Because each critic satisfies $\operatorname{Var}\big[Q_{\theta_{k}}(s_{t},a_{t})\big]\leq\sigma^{2}$ for all $k$, the expected standard deviation is bounded, and therefore
$\lim_{N\to\infty}\mathbb{E}\bigg[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\bigg]\leq\beta_{\min}\cdot\sigma.$ (18)
In particular, when $\beta_{\min}\to 0$,
$\lim_{N\to\infty}\mathbb{E}\bigg[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\bigg]\leq 0,$ (19)
which implies
$\mathbb{E}\bigg[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\bigg]\xrightarrow{N\to\infty}0.$ (20)
Hence, the ensemble estimator $Q_{E}(s_{t},a_{t})$ converges in probability to the true value function $Q^{\ast}(s_{t},a_{t})$ as the number of training iterations grows to $\infty$ and $\beta_{\min}$ tends to zero.
Appendix B Experimental Details
B.1 Environments
The MuJoCo environments used in our experiments (Section 5) follow the same configuration as QARL (Reddi et al., 2024), with the magnitude of the adversarial force tailored to each environment. To avoid catastrophic policy collapse, the maximum adversarial force for each environment is explicitly capped. All environment-specific parameters are listed in Table 3. Note that the adversary action spaces reported in Table 3 do not apply to CAT (Sheng et al., 2022) and MixedNE-LD (Kamalaruban et al., 2020), as in these methods the protagonist and adversary share the same action space.
B.2 Hyperparameter Setting
   Hyperparameter    Value
   Shared (ALL)
    # iterations NN    200
    # episodes per agent per iteration    5
    # evaluation rollouts per iteration    10
    # hidden layers    3
    # hidden units per layer    256
    discount factor    0.99
    horizon    500
   UACER
    # ensemble Critic networks $K$    5
    $\beta_{0}$    0.85
    $\beta_{min}$    0.15
    decay speed $\lambda$    3
   Shared (SAC and SAC MixedNE-LD)
    nonlinearity    ReLU
    critic optimiser    Adam
    critic learning rate    $3\times 10^{-4}$
    actor learning rate    $1\times 10^{-4}$
    initial replay memory size    $3\times 10^{3}$
    max replay memory size    $1\times 10^{6}$
    warmup transitions    $5\times 10^{3}$
    batch size    256
    target smoothing coefficient ($\tau$)    $5\times 10^{-3}$
    target update interval    1
    policy log std bounds    [20,2][-20,2]
    initial temperature    $5\times 10^{-3}$
    temperature learning rate    $3\times 10^{-4}$
    target entropy    $\textrm{dim}(\mathcal{A})$
   SAC
    actor optimiser    Adam
   SAC MixedNE-LD
    adversary influence $\sigma$    0.1
    actor optimiser    SGLD
    thermal noise ($\sigma_{t}$)    $10^{-3}\times(1-5\times 10^{-5})^{t}$
    RMSProp parameter $\alpha$    0.999
    RMSProp parameter $\epsilon$    $10^{-8}$
   CAT
    curriculum start iteration    $0.2\times\text{\# iterations}$
    curriculum end iteration    $0.8\times\text{\# iterations}$
   CAT Adversary
    gradient descent learning rate    3
    gradient descent step limit    25
    gradient descent convergence threshold $\epsilon$    $10^{-3}$
    disturbance $L^{p}$-norm    2
   QARL
    initial gamma distribution concentration $k_{\text{initial}}$    50
    target gamma distribution concentration $k_{\text{target}}$    1
    fixed gamma distribution rate    1000
    $D_{\text{KL}}$ constraint ($\epsilon$)    0.5
    # rollouts needed for update    30
Table 4: Algorithm hyperparameters.
To ensure fair and valid comparisons, we strictly follow the experimental protocol established by QARL (Reddi et al., 2024), maintaining consistent shared hyperparameter settings across all methods. This practice ensures that performance differences can be attributed to algorithmic distinctions rather than variations in parameter tuning. The hyperparameters used for UACER and each baseline method are provided in Table 4.
B.3 Hardware and Software
All experiments were run on a workstation equipped with an NVIDIA TITAN Xp GPU (12 GB), 15 GB of system memory, and an Intel® Xeon® E5-2680 v4 CPU. The implementations of the algorithms are based on the publicly available codebase of Reddi et al. (2024), which itself is built on the MushroomRL library (D’Eramo et al., 2021). The same library is also used to implement the agents and to construct the adversarial environment wrappers.